OpenStudy (anonymous):

3d + c = 7 2d - c = 6 solve for d and c

OpenStudy (anonymous):

add the two equations together lined up as you have it

OpenStudy (anonymous):

\[3d+2d=5d\] \[-c+c=0\] \[7+6=13\]

OpenStudy (anonymous):

so you get \[5d+0=13\] or just \[5d=13\] then to get d divide by 5 and have \[d=\frac{13}{5}\]

OpenStudy (anonymous):

now replace d by \[\frac{13}{5}\] ini either equation and solve for c via \[3\times \frac{13}{5}+c=7\] or \[2\times \frac{13}{5}-c=6\]

OpenStudy (anonymous):

i will pick the first one \[\frac{39}{5}+c=7\] \[c=7-\frac{39}{5}\] \[c=-\frac{4}{5}\]

OpenStudy (anonymous):

i know now thanks ")

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

its actualy so simple lol...eh guess im just tired